OpenStudy (anonymous):

Explain the steps in solving the following absolute value equation. 2|x-2|=12. plz HELP

OpenStudy (anonymous):

@satellite73 plz help

OpenStudy (anonymous):

@ikram002p help?

OpenStudy (anonymous):

is the answer x= 4 and 8? @ikram002p

OpenStudy (ikram002p):

2|x-2|=12. |x-2|=6 x=\( \left(\begin{matrix} x-2, x >=2 \\ -x+2 , x <-2\end{matrix}\right)\). case 1 :- x-2=6 x=8 case 2 2-x=6 x=4

OpenStudy (ikram002p):

yes ur correct its 4,8

OpenStudy (anonymous):

yay ok i got that part right. whats next?

OpenStudy (anonymous):

is that it?

OpenStudy (ikram002p):

done nothing next :)

OpenStudy (ikram002p):

yep

OpenStudy (anonymous):

thank you!

OpenStudy (ikram002p):

np :)